# Pushing the Limits of Distillation-Based Class-Incremental Learning via Lightweight Plugins

### v2

@@ 5.2 Effectiveness on Distillation-Based CIL
In this section, we evaluate the performance improvement brought by the proposed DLC framework across different distillation-based CIL objectives, including iCaRL [23], WA [38], BiC [31], DAKD [12], and MKD [20]. Additionally, a simple herding-based Replay method [23] is included to examine DLC’s contribution in scenarios relying solely on the rehearsal without distillation. We report the $\overline{A}$ and $A_{T}$ of these established methods with and without DLC across three datasets under varying task configurations. In all experiments, the replay buffer size is fixed at 2,000 exemplars. DLC expands only one LoRA plugin per task, applied to the last convolutional layer of the feature extractor. We also report the total parameters required by DLC and the size of the entire feature extractor for comparison. The results are summarized in Table 1.
The results show that DLC significantly improves all classical replay or distillation-based methods with only one LoRA plugin per task. Moreover, the parameter overhead introduced by DLC is negligible. DLC achieves up to 8% accuracy improvement while expanding only 4.8% to 32% parameters equivalent to the feature extractor, highlighting its practical value in CIL. Fig. 3 shows the confusion matrix heatmap of iCaRL and iCaRL-DLC as an example. DLC increases the thermal intensity along the diagonal for most classes, demonstrating that it could enhance the model’s accuracy across most tasks.
In addition, we compare the final accuracy of DLC-enhanced iCaRL, WA, and BiC with and without the weighting unit, as shown in Fig. 4. Results indicate that the proposed weighting unit contributes a certain accuracy gain.
@@ 5.3 Comparison under Fixed Memory Budget
Table 2: Performance comparison on CIFAR-100 with aligned memory cost. ‘#P’ represents the number of parameters (million). $\varepsilon$’ denotes the number of exemplars, and ‘MS’ denotes the memory size (MB). All methods use ResNet-32 as the backbone.
@@
Prior research [42] has demonstrated that model-level expansion-based methods achieve state-of-the-art performance across all standard CIL benchmarks. However, these approaches overlook the substantial memory overhead induced by dynamic architecture growth. Such methods implicitly introduce additional memory budgets, i.e., a model buffer for retaining parameters of the expanded model, which grants them a considerable and unaccounted advantage over other methods that do not store expansion models or adopt more lightweight expansion strategies.
Following the protocol established in [43], this section presents a memory-aligned comparison of CIL baselines. Under a fixed 23.5 MB memory budget, the memory not consumed by a method’s parameters is allocated to replay exemplars. To balance this, methods with fewer parameters store more exemplars, and vice versa. This ensures an equitable evaluation. In addition to the baselines used in Section 5.2, we add several state-of-the-art approaches for comprehensive comparison, including FOSTER [29], DER [33], MEMO [43], and TagFex [39]. Among them, DER, MEMO, and TagFex have incorporated the weight alignment from WA in their implementations. As summarized in Table 2, deploying DLC to the distillation-based method WA under the memory-aligned setting yields the highest classification accuracy.
5.4 Compatibility with other Plug-and-Play Enhancements
Figure 5: Comparison of $A_{T}$ with/without DLC on CIFAR-100 when enhanced with AutoAugment (AA) and $\mathcal{L}_{\text{inter-intra}}$.
In addition to DLC, other plug-and-play strategies have been proposed to enhance CIL methods. For example, the data augmentation technique AutoAugment [6] has been employed in the CIFAR-100 dataset by [29, 39]. Meanwhile, two additional training losses $\mathcal{L}_{\text{inter}}$ and $\mathcal{L}_{\text{intra}}$ specifically designed for logit-based distillation methods have been introduced in [10]. This subsection evaluates the compatibility of DLC with such enhanced strategies. Specifically, we compare the accuracy of methods using AutoAugment and the loss $\mathcal{L}_{\text{inter-intra}}$ with and without DLC. For experiments with $\mathcal{L}_{\text{inter-intra}}$, we follow the original implementation and use ResNet-18 as the backbone. As shown in the Fig. 5, DLC consistently brings significant improvements, demonstrating its compatibility with other plug-and-play enhancements.
5.5 Additional Computational Overhead from DLC
This section investigates the additional inference cost introduced by deploying DLC. Using iCaRL as a representative baseline, we evaluate two deployment variants: iCaRL equipped with one LoRA plugin per task, and iCaRL equipped with both the LoRA plugins and the weighting unit. Concretely, on ImageNet-100, we report the total FLOPs and the average inference latency per sample (ms) measured on the same test set. The results are shown in Fig. 6. Overall, neither FLOPs nor latency increases substantially as the number of tasks grows. Instead, both grow mildly as the task sequence extends. Even under long sequences such as B5 Inc5, the increases in FLOPs and inference time are only about 4%. These results indicate that DLC provides a lightweight and computation-friendly parameter extension for enhancing distillation-based CIL.
Figure 6: Comparison of inference FLOPs and per-sample inference latency with and w/o DLC under different task settings.
@@ 6 Conclusion
In this work, we address a challenge in distillation-based class-incremental learning: improving the model performance with only a small amount of task-specific auxiliary parameters, without breaking the single-model distillation setting. We propose DLC, a plug-and-play plugin extension framework that augments replay and distillation-based baselines with task-specific LoRA-style residual plugins. DLC trains each plugin in a decoupled manner and then freezes it, which both avoids interference from distillation objectives and preserves the acquired task knowledge. We further introduce a lightweight weighting unit to down-weight irrelevant plugin features at inference. Extensive experiments demonstrate that DLC consistently improves replay and distillation-based CIL baselines with few parameter overhead. Future work will explore more expressive weighting designs for stronger feature integration, as well as earlier-layer plugin deployment to further improve adaptability.
Appendix
Appendix A Dataset Description
CIFAR-100 CIFAR-100 (Canadian Institute for Advanced Research 100-class dataset) is a widely used standard benchmark dataset in the field of computer vision, particularly for image classification tasks. It is an expanded version of the CIFAR-10 dataset, curated by the Canadian Institute for Advanced Research. A defining feature of this dataset is its fine-grained categorical structure. It contains a total of 100 distinct fine-grained classes (fine labels), organized hierarchically under 20 broader superclasses (coarse labels), with each superclass encompassing precisely 5 fine-grained classes. This hierarchical organization provides a valuable foundation for researching multi-level classification. The dataset comprises 60,000 tiny color images, each with a fixed resolution of 32x32 pixels. This compact size makes the dataset manageable for rapid experimentation and algorithm iteration, while simultaneously presenting significant challenges for model recognition capabilities due to the extreme low resolution. The 60,000 images are explicitly divided into 50,000 training images and 10,000 test images, meaning each fine-grained class includes 500 training images and 100 test images. CIFAR-100 offers broad and diverse category coverage, encompassing common objects (various vehicles, furniture, appliances), animals (mammals, fish, insects), plants (flowers, trees), and natural scenes (mountains, oceans), among others. Its extensive number of classes combined with its moderate scale make it an ideal testbed for evaluating and comparing the performance and generalization capabilities of image classification models—especially Convolutional Neural Networks (CNNs)—as well as for exploring advanced tasks like few-shot learning and transfer learning.
ImageNet-100 ImageNet-100 is a commonly used subset of the well-known large-scale image dataset, ImageNet. Its purpose is to significantly reduce computational resource requirements and training time while preserving much of ImageNet’s richness and challenge, making it more suitable for rapid prototyping, algorithm exploration, and research in resource-constrained environments. As the name implies, ImageNet-100 consists of 100 classes (synsets) selected from the original ImageNet dataset. The selection of these classes is typically not entirely random; it often follows specific strategies (e.g., choosing semantically distinct classes or classes with relatively higher sample counts) to ensure the subset remains representative and diverse. The dataset contains approximately 100,000 high-resolution color images curated from these 100 classes. This represents a substantial reduction in scale compared to the millions of images in the full ImageNet. The standard partition of the dataset usually includes a training set and a validation set. The training set contains roughly 1,000 images per class, totaling approximately 100,000 images. The validation set contains about 50 images per class, totaling around 5,000 images. The images themselves are typically of variable sizes (significantly larger than CIFAR’s 32x32 pixels) and require preprocessing (like cropping and resizing) before being fed into models. ImageNet-100 inherits core characteristics from its parent dataset: rich and realistic image content, high semantic distinction between classes, significant intra-class variation (the same object under different viewpoints, lighting, occlusion), and inter-class similarities. It is widely used for training and evaluating image classification models, as well as for transfer learning tasks. Compared to smaller datasets like CIFAR, models pre-trained on ImageNet-100 generally learn more generalizable visual features suitable for fine-tuning on downstream tasks. Compared to the full ImageNet, it offers faster iteration cycles and lower hardware barriers to entry.
TinyImageNet-200 TinyImageNet-200 is a downsized variant of the ImageNet dataset, specifically created for educational and experimental purposes in deep learning and computer vision research. It features a curated subset of 200 distinct object categories (synsets) carefully selected from the full ImageNet hierarchy. The dataset contains roughly 100,000 high-resolution color images – significantly fewer than its parent dataset but substantially more than smaller benchmarks like CIFAR. Each image in Tiny ImageNet has been consistently downscaled and cropped to a fixed resolution of 64x64 pixels, maintaining a usable size while greatly reducing computational demands compared to processing full-sized ImageNet images. The dataset comes with a predefined partition: a training set containing 500 images per class, a separate validation set with 50 images per class, and a test set also containing 50 images per class, totaling approximately 100,000 training images, 10,000 validation images, and 10,000 test images. Alongside the images, TinyImageNet provides essential metadata files including a wnids.txt file listing the WordNet IDs for all 200 classes, a words.txt file mapping these IDs to human-readable labels, and a structured val_annotations.txt file listing the ground-truth classes for validation images. Its manageable scale, balanced class representation, and standardized structure make Tiny ImageNet-200 a popular choice for academic courses exploring deep learning architectures (particularly CNNs and Vision Transformers), as well as for research prototyping on tasks such as image classification, transfer learning, and model robustness evaluation, especially where computational resources are constrained but larger challenge than CIFAR is needed.
Appendix B Feature Drift Bounds under Logit Distillation and Herding Replay
Before applying methods like LoRA to continually updated models, it is imperative to guarantee that the feature extractor’s output deviation for previous tasks remains bounded before and after training. Otherwise, if catastrophic forgetting causes a significant shift in outputs, the plugin modules would be unable to extract meaningful residuals.
Setting.
Consider a sequence of tasks $D_{1},D_{2},\dots,D_{t}$. At task $t$, the model is trained on
$D_{\mathrm{train}}=D_{t}\cup D_{\exp},$ (13)
where $D_{\exp}$ is the class-balanced exemplar set from previous tasks based on the herding replay strategy in iCaRL. Let $\theta$ denote the parameters of a convolutional feature extractor $\phi(\boldsymbol{x})$, and let $h_{\ell}(x;\theta)\in\mathbb{R}^{d_{\ell}}$ denote the output of layer $\ell$ in $\phi(\boldsymbol{x})$. We denote the parameters of $\phi(\boldsymbol{x})$ as $\theta_{t-1}$ and $\theta_{t}$ for before and after training on task $t$.
We adopt the logit-based distillation approach as a representative case for analysis. To constrain forgetting, the logits are distilled from the teacher to the student. The logits of teacher and student are denoted as $\hat{z}(x)$ and $z(x)$. For temperature $T>0$, define the softened probabilities
$\hat{q}^{T}(x)=\mathrm{softmax}\!\big(\hat{z}(x)/T\big),$ (14)
$q^{T}(x)=\mathrm{softmax}\!\big(z(x)/T\big).$ (15)
The cross-entropy-based distillation loss is
$\mathcal{L}_{\mathrm{KD}}(x)=-\sum_{i=1}^{C}\hat{q}^{T}_{i}(x)\log q^{T}_{i}(x)$ (16)
$=H\!\big(\hat{q}^{T}(x)\big)+\mathrm{KL}\!\big(\hat{q}^{T}(x)\|q^{T}(x)\big).$ (17)
Where $H(\hat{q}^{T}(x))$ denotes the Shannon entropy of the teacher distribution,
$H(\hat{q}^{T}(x)):=-\sum_{i=1}^{C}\hat{q}^{T}_{i}(x)\log\hat{q}^{T}_{i}(x),$
so the KD loss decomposes into the teacher’s entropy and the KL divergence between the teacher and the student.
The empirical distillation loss over $D_{\exp}$ is
$\bar{L}_{\mathrm{KD}}=\frac{1}{|D_{\exp}|}\sum_{x\in D_{\exp}}\mathcal{L}_{\mathrm{KD}}(x).$ (18)
It serves as a measurable proxy for how closely the student imitates the teacher on the subset of old-task samples preserved through herding replay.
Our goal is to study how much the intermediate-layer features drift after learning on a new task. For any layer $\ell$, define the drift as:
$\Delta h_{\ell}(x)=h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1}).$ (19)
which quantifies the amount by which the feature extractor changes its representation of an old-task sample. Bounding this is crucial for ensuring that LoRA plugins trained on earlier feature distributions remain effective after the model is updated on new tasks.
In addition, to compare logits in a way that reflects changes in their meaningful structure rather than arbitrary additive shifts, we use centered logits:
$\bar{z}(x)=z(x)-\tfrac{1}{C}\sum_{k=1}^{C}z_{k}(x)\mathbf{1},$ (20)
$\bar{\hat{z}}(x)=\hat{z}(x)-\tfrac{1}{C}\sum_{k=1}^{C}\hat{z}_{k}(x)\mathbf{1}.$ (21)
Centered logit is necessary because logits are defined only up to a constant offset: adding a constant to all logit entries leaves the softmax distribution unchanged. By removing this offset, we ensure that the norms of logit differences reflect genuine changes in class-relative predictions rather than uninformative bias shifts. This step allows the later Lipschitz bounds to connect logit discrepancies with feature discrepancies in a well-defined way.
Assumptions
A1 (Non-degenerate teacher distribution). There exists $\delta\in(0,1/C]$ that satisfies:
$\hat{q}^{T}_{i}(x)\geq\delta\quad\text{for all }x\in D_{\exp}\text{ and }i.$ (22)
This condition prevents extreme logit sharpness and ensures the log-softmax function is well-conditioned. The assumption is directly satisfied in standard knowledge distillation settings. In particular, the use of a temperature parameter $T>1$ produces softened teacher probabilities,
$\hat{q}^{T}_{i}(x)=\frac{\exp(\hat{z}_{i}(x)/T)}{\sum_{j}\exp(\hat{z}_{j}(x)/T)},$
which guarantees that no class probability collapses to zero. Alternatively, label smoothing also enforces a uniform lower bound on class probabilities. Thus, A1 follows from standard KD practice and is not a restrictive condition.
A2 (Local bi-Lipschitz continuity). Let $\Phi_{\ell}$ denote the mapping from the $\ell$-th layer feature to the centered logits $\bar{z}$ through subsequent layers. There exist constants $0<\mu_{\ell}\leq L_{\ell}<\infty$ such that
$\mu_{\ell}\|u-v\|_{2}\leq\|\Phi_{\ell}(u)-\Phi_{\ell}(v)\|_{2}\leq L_{\ell}\|u-v\|_{2},$ (23)
for all $u,v$ within a neighborhood of past-task features in $D_{\exp}$. This ensures logits and features change proportionally, which is well supported by the structure of modern neural networks. Convolutional layers, affine transformations, and common nonlinearities (ReLU, SiLU, GELU) are all Lipschitz continuous. Since compositions of Lipschitz functions remain Lipschitz, the entire feature-to-logit map possesses an upper Lipschitz constant. The lower Lipschitz constant reflects local non-collapse of the representation: empirically, trained deep networks do not map distinct inputs from the same manifold region to identical logits. Similar assumptions have been used in analyses of embedding stability and invertibility of deep models. Therefore, A2 expresses a standard geometric regularity of neural networks rather than a special structural requirement.
A3 (Empirical KD value). At the end of training, the empirical KD loss over $D_{\exp}$ equals $\bar{L}_{\mathrm{KD}}$. This is not an additional assumption but rather a definition of the observable optimization outcome. The analysis treats $\bar{L}_{\mathrm{KD}}$ as a known scalar summarizing how well the student matches the teacher on replay samples.
A4 (Generalization from herded exemplars). The replay buffer $D_{\exp}$ maintains an equal number of samples per class, and $\mathcal{L}_{\mathrm{KD}}(x)\in[0,\log C]$ is bounded, so with probability at least $1-\delta_{0}$, we have:
$\big|\mathbb{E}_{x\sim\mathcal{D}_{<t}}[\mathcal{L}_{\mathrm{KD}}(x)]-\bar{L}_{\mathrm{KD}}\big|\leq\epsilon_{\mathrm{gen}},$ (24)
$\epsilon_{\mathrm{gen}}=O\!\Big(\sqrt{\tfrac{\log(1/\delta_{0})}{|D_{\exp}|}}\Big),$ (25)
where $\mathcal{D}_{<t}$ is the joint distribution of all previous tasks. This assumption holds because the following properties guarantee that the empirical distillation loss concentrates around its population expectation. The boundedness of the distillation loss,
$0\leq\mathcal{L}_{\mathrm{KD}}(x)\leq\log C,$
implies that empirical averages of $\mathcal{L}_{\mathrm{KD}}$ satisfy standard concentration inequalities. Moreover, the herding procedure constructs $D_{\exp}$ as a class-balanced set that approximates class means in feature space, making it an approximately representative subset of the past-task distribution $\mathcal{D}_{<t}$. As a result, the empirical quantity
$\bar{L}_{\mathrm{KD}}=\frac{1}{|D_{\exp}|}\sum_{x\in D_{\exp}}\mathcal{L}_{\mathrm{KD}}(x)$
forms a low-variance estimator of the population expectation. Applying Hoeffding-type bounds for bounded losses yields, with probability at least $1-\delta_{0}$,
$\big|\mathbb{E}_{x\sim\mathcal{D}_{<t}}[\mathcal{L}_{\mathrm{KD}}(x)]-\bar{L}_{\mathrm{KD}}\big|\;\leq\;O\!\left(\sqrt{\tfrac{\log(1/\delta_{0})}{|D_{\exp}|}}\right),$
Auxiliary Lemmas
Lemma 1 (KD controls probability discrepancy). For any $x$, by Pinsker’s inequality:
$\|q^{T}(x)-\hat{q}^{T}(x)\|_{1}\leq\sqrt{2\,\mathrm{KL}(\hat{q}^{T}(x)\|q^{T}(x))}\leq\sqrt{2(\mathcal{L}_{\mathrm{KD}}(x)-H(\hat{q}^{T}(x)))}.$ (26)
Averaging this equation over $D_{\exp}$ gives
$\frac{1}{|D_{\exp}|}\sum_{x\in D_{\exp}}\|q^{T}(x)-\hat{q}^{T}(x)\|_{1}\leq\sqrt{2\Big(\bar{L}_{\mathrm{KD}}-\tfrac{1}{|D_{\exp}|}\sum_{x}H(\hat{q}^{T}(x))\Big)}.$
Lemma 2 (Probability discrepancy controls logit discrepancy). Softmax relates logits and probabilities by:
$(z_{i}-z_{j})=T\log(\frac{q^{T}_{i}}{q^{T}_{j}}).$ (27)
So under A1, for any $i\neq j$, Because of Lipschitz continuity of $\log p$ over a bounded domain, differences in log-probabilities are linearly bounded by differences in probabilities:
$|\log q^{T}_{i}-\log\hat{q}^{T}_{i}|\leq\frac{T}{\delta}(|q^{T}_{i}-\hat{q}^{T}_{i}|).$ (28)
So it has:
$|(z_{i}-z_{j})-(\hat{z}_{i}-\hat{z}_{j})|=T\Big|\log\frac{q^{T}_{i}}{q^{T}_{j}}-\log\frac{\hat{q}^{T}_{i}}{\hat{q}^{T}_{j}}\Big|$ (29)
$\leq\frac{T}{\delta}(|q^{T}_{i}-\hat{q}^{T}_{i}|+|q^{T}_{j}-\hat{q}^{T}_{j}|).$ (30)
Therefore,
$\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}\leq\frac{T}{\delta}\|q^{T}(x)-\hat{q}^{T}(x)\|_{1}.$ (31)
Averaging Eq.(31) over $D_{\exp}$ and combining it with Eq. (B) in Lemma 1 yields:
$\frac{1}{|D_{\exp}|}\sum_{x\in D_{\exp}}\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}\leq\frac{T}{\delta}\sqrt{2\Big(\bar{L}_{\mathrm{KD}}-\tfrac{1}{|D_{\exp}|}\sum_{x}H(\hat{q}^{T}(x))\Big)}.$ (32)
Lemma 3. Under A2, for any $x$, it has:
$\mu_{\ell}\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\|_{2}\leq\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}$ (33)
$\leq L_{\ell}\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\|_{2}.$ (34)
Main Conclusions
Theorem 1 (Empirical bound on feature drift). For any layer $\ell$,
$\frac{1}{|D_{\exp}|}\sum_{x\in D_{\exp}}\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\|_{2}\leq\frac{T}{\mu_{\ell}\delta}\sqrt{2\Big(\bar{L}_{\mathrm{KD}}-\tfrac{1}{|D_{\exp}|}\sum_{x}H(\hat{q}^{T}(x))\Big)}.$ (35)
Proof. This conclusion can be straightforwardly proved by combining the Eq.(33) in Lemma 3 (first averaging it over $D_{\exp}$ ) with the Eq.(32) of Lemma 2.
Theorem 2 (Distributional bound over past-task distribution). Under A1–A4, with probability at least $1-\delta_{0}$,
$\mathbb{E}_{x\sim\mathcal{D}_{<t}}\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\|_{2}\leq\frac{T}{\mu_{\ell}\delta}\sqrt{2\Big(\bar{L}_{\mathrm{KD}}-\mathbb{E}_{x\sim\mathcal{D}_{<t}}[H(\hat{q}^{T}(x))]+\epsilon_{\mathrm{gen}}\Big)}.$ (36)
Proof. The expectation form of Lemma 2 states that
$\ \mathbb{E}_{x\sim\mathcal{D}_{<t}}\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}\;\leq\;\frac{T}{\delta}\sqrt{2\Big(\mathbb{E}_{x\sim\mathcal{D}_{<t}}\mathcal{L}_{\mathrm{KD}}(x)-\mathbb{E}_{x\sim\mathcal{D}_{<t}}H(\hat{q}^{T}(x))\Big)}.$ (37)
A4 ensures that the empirical distillation loss generalizes to the past-task distribution:
$\Big|\mathbb{E}_{x\sim\mathcal{D}_{<t}}\mathcal{L}_{\mathrm{KD}}(x)-\bar{L}_{\mathrm{KD}}\Big|\;\leq\;\epsilon_{\mathrm{gen}}.$ (38)
Hence,
$\mathbb{E}_{x\sim\mathcal{D}_{<t}}\mathcal{L}_{\mathrm{KD}}(x)\;\leq\;\bar{L}_{\mathrm{KD}}+\epsilon_{\mathrm{gen}}.$ (39)
Substituting this into Eq.(37) yields:
$\mathbb{E}_{x\sim\mathcal{D}_{<t}}\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}\;\leq\;\frac{T}{\delta}\sqrt{2\Big(\bar{L}_{\mathrm{KD}}-\mathbb{E}_{x\sim\mathcal{D}_{<t}}H(\hat{q}^{T}(x))+\epsilon_{\mathrm{gen}}\Big)}.$ (40)
Lemma 3 states that for every $x$,
$\mu_{\ell}\big\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\big\|_{2}\;\leq\;\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}.$ (41)
Taking expectation on both sides gives:
$\mathbb{E}_{x\sim\mathcal{D}_{<t}}\big\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\big\|_{2}\;\leq\;\frac{1}{\mu_{\ell}}\mathbb{E}_{x\sim\mathcal{D}_{<t}}\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}.$ (42)
Then the conclusion to be proved can be obtained by combining Eq.(40) with Eq.(42).
Theorems 1–2 imply
$\mathbb{E}\|\Delta h_{\ell}(x)\|\leq\underbrace{\tfrac{T}{\mu_{\ell}\delta}}_{\text{constant }K_{\ell}}\times\underbrace{\sqrt{2(\Delta_{\mathrm{KD}}+\epsilon_{\mathrm{gen}})}}_{\text{variable }\Gamma_{t}},$ (43)
where $\Delta_{\mathrm{KD}}=\mathbb{E}[\mathcal{L}_{\mathrm{KD}}(x)-H(\hat{q}^{T}(x))]$ reflects the effectiveness of distillation on the data of previous tasks. A smaller $\Delta_{\mathrm{KD}}$ represents the predictive behavior of the student model that aligns with that of the teacher model on these tasks. $\epsilon_{\mathrm{gen}}=O(\sqrt{\log(1/\delta_{0})/|D_{\exp}|})$ decreases as more exemplars are stored. The constant $\tfrac{T}{\mu_{\ell}\delta}$ depends on network smoothness and temperature parameters. The $K_{\ell}$ captures the conditioning of the feature-to-logit map and the softness of the teacher distribution, while the $\Gamma_{t}$ measures how much logit drift the distillation and replay objective permits.
This bound shows that when distillation is optimized and the exemplars approximate the old-task distribution well, the deep feature representations of past data remain stable even as the model learns new tasks. Therefore, even though $\phi(\boldsymbol{x})$ continuously updates its parameters, the output of each layer for previous tasks remains relatively stable. This allows the plugins trained on earlier tasks to stay effective, thereby enabling the deployment of DLC in replay- and distillation-based CIL methods.

### v3

@@ 5.2 Effectiveness on Distillation-Based CL
In this section, we evaluate the performance improvement brought by the proposed DLC framework across different distillation-based CL objectives, including iCaRL [27], WA [43], BiC [36], DAKD [13], and MKD [24]. Additionally, a simple herding-based Replay method [27] is included to examine DLC’s contribution in scenarios relying solely on the rehearsal without distillation. We report the $\overline{A}$ and $A_{T}$ of these established methods with and without DLC across three datasets under varying configurations. In all experiments, the replay buffer size is fixed at 2,000 exemplars. DLC expands only one LoRA plugin per stage, i.e., each plugin consists of a single pair of low-rank matrices $\mathbf{A}$ and $\mathbf{B}$. We also report the total parameters required by DLC and the size of the entire feature extractor for comparison. The results are summarized in Table 1. The results show that DLC significantly improves all distillation-based methods. Moreover, the parameter overhead introduced by DLC is negligible. DLC achieves up to 8% accuracy improvement while expanding only 4.8% to 32% parameters equivalent to the feature extractor, highlighting its practical value in CL. Fig. 4 shows the confusion matrix heatmap of iCaRL and iCaRL-DLC as an example. DLC increases the thermal intensity along the diagonal for most classes, demonstrating that it could enhance the model’s accuracy across most classes.
@@
Figure 6: Comparison of $A_{T}$ with plugins deployed at different positions (FL/ML/CPL) on CIFAR-100.
In addition to DLC, other plug-and-play enhancement strategies have been proposed to enhance CL methods. For example, some works employ the data augmentation strategy AutoAugment for continual learning, particularly on the CIFAR-100 dataset [33, 44]. Several plug-and-play loss functions have also been introduced, including $\mathcal{L}_{\text{inter}}$ and $\mathcal{L}_{\text{intra}}$ for distillation methods [8], as well as C-flat [18]. We evaluate the compatibility of DLC with such enhanced strategies. Specifically, we compare the $A_{T}$ of baseline methods using AutoAugment, the loss $\mathcal{L}_{\text{inter-intra}}$, and C-flat with and without DLC. For experiments with $\mathcal{L}_{\text{inter-intra}}$, we follow the original implementation and use ResNet-18 as the backbone. As shown in the Fig. 3, DLC consistently brings significant improvements, demonstrating its compatibility with other plug-and-play enhancements.
Moreover, we conduct two ablation studies. First, we evaluate the effect of the proposed weighting unit by comparing the final accuracy of DLC-enhanced iCaRL, WA, and BiC with and without it, as shown in Fig. 5. Results indicate that the weighting unit contributes to improved performance. Second, we investigate the importance of classifier-proximal deployment by placing the plugin at different depths of the feature extractor, including the first layer (FL), middle layer (ML), and the classifier-proximal layer (CPL). As shown in Fig. 6, only classifier-proximal deployment improves performance, whereas earlier deployment causes increasingly severe degradation, indicating poorer distillation compatibility.
Figure 7: Comparison of $A_{T}$ with/without DLC under challenging continual learning settings, including long-sequence, few-shot, and class-imbalanced scenarios.
5.3 Evaluation under Challenging CL Settings
Distillation-based CL methods are sometimes deployed in extreme scenarios, such as long learning sequences, few-shot, and class-imbalanced data. We therefore further evaluate DLC under these settings. For the long-sequence setting, we use Tiny-ImageNet with the B5Inc5 protocol, resulting in 40 stages. For the few-shot setting, we follow the protocol in [16] on CIFAR with B60Inc5, where each class contains only 5 samples after the second stage. For the class-imbalance setting, following Liu et al. [21], we adopt the CIFAR-100 B50Inc5 protocol with an imbalance ratio $\rho=0.01$ (i.e., head-to-tail ratio takes 100). The results are shown in Fig. 7. DLC consistently improves distillation-based CL methods even under these challenging settings.
Figure 8: Comparison of inference FLOPs and per-sample inference latency with and w/o DLC.
@@ 5.4 Comparison under Fixed Memory Budget
Table 2: Performance comparison on CIFAR-100 with aligned memory cost. ‘#P’ represents the number of model parameters (million). $\varepsilon$’ denotes the number of exemplars, and ‘MS’ denotes the total memory size (MB).
@@
Prior studies have shown that model-level expansion methods can achieve state-of-the-art performance in continual learning, but their gains often rely on substantial and uncounted memory overhead for storing expanded model parameters, making direct comparison with distillation-based methods unfair. To further demonstrate the parameter efficiency and accuracy of DLC under a fair budget constraint, we follow the protocol established in [48] and present a memory-aligned comparison of CL baselines. Under a fixed 23.5 MB memory budget, the memory not consumed by a method’s parameters is allocated to replay exemplars. To balance this, methods with fewer parameters store more exemplars, and vice versa. This ensures an equitable evaluation. In addition to the baselines used in Section 5.2, we add several state-of-the-art approaches for comprehensive comparison, including FOSTER [33], DER [38], MEMO [48], and TagFex [44]. Among them, DER, MEMO, and TagFex have incorporated the weight alignment from WA in their implementations. All methods use the same ResNet backbone. As summarized in Table 2, WA equipped with DLC achieves the highest classification accuracy under the memory-aligned setting. This highlights that DLC delivers a superior memory-accuracy trade-off.
5.5 Additional Computational Overhead
This section investigates the additional inference cost introduced by deploying DLC. Using iCaRL as a representative baseline, we evaluate two deployment variants: iCaRL equipped with only DLC plugins, and iCaRL equipped with both the plugins and the weighting unit. Concretely, on ImageNet-100, we report the total FLOPs and the average inference latency per sample (ms) measured on the same test set. The results are shown in Fig. 8. Overall, neither FLOPs nor latency increases substantially as the number of stages grows. Instead, both grow mildly as the stage sequence extends. Even under long sequences such as B5 Inc5, the increases in FLOPs and inference time are only about 4%. These results indicate that DLC provides a lightweight and computation-friendly parameter extension for enhancing distillation-based CL.
@@ 6 Conclusion
In this work, we address a challenge in distillation-based continual learning: improving the model performance with only a small amount of auxiliary parameters, without breaking the single-model distillation setting. We propose DLC, a plug-and-play plugin extension framework that augments replay and distillation-based baselines with LoRA-style residual plugins. DLC deploys each plugin at the classifier-proximal layer and trains it in a decoupled manner, allowing residual correction to directly refine features while remaining compatible with distillation. We further introduce a lightweight weighting unit to down-weight irrelevant features in inference. Extensive experiments demonstrate that DLC consistently improves distillation-based CL baselines with few parameter overhead. Future work will explore more expressive weighting designs for stronger feature integration, as well as earlier-layer plugin deployment to further improve adaptability.